# Fabrication and characterization of optical filters from polymeric aerogels loaded with diamond scattering particles

### v2

@@
We have developed a suite of infrared-blocking filters made by embedding diamond scattering particles in a polyimide aerogel substrate. We demonstrate the ability to tune the spectral performance of the filters based on both the composition of the base aerogel material and the properties of the scattering particles. We summarize the fabrication, optical modeling, and characterization of these filters. We investigate two polyimide base aerogel formulations and the effects of loading them with diamond scattering particles of varying sizes and relative densities. We describe a model for the filters’ behavior using a combination of Maxwell Garnett and Mie Scattering techniques. We present optical characterization results for diamond-loaded aerogel filters with cutoff frequencies (50% transmittance) ranging between 2.5 and 15 THz, and confirm that the measured spectral performance is in agreement with our optical models. We also measure the filters’ refractive index in the microwave and report findings in agreement with Maxwell Garnett model predictions (typically $n<1.08$).
@@ 1 Introduction
Due to the high porosity of aerogels, they can serve as a host system to other materials without significantly compromising their performance as lightweight, low density, low permittivity materials. The available free volume of the aerogel material allows for the incorporation of particles of higher index material during fabrication (see Section 2.1). The resulting composite material consists of a network of high index particles held in suspension by the low index aerogel[24]. Such particle-loaded polymeric aerogel composites can subsequently be used as optical low-pass filters in the far-infrared through microwave regimes. Incoming light in the signal band passes through the very low-reflectance and low-loss base medium while unwanted higher frequency light is scattered and reflected by the composite aerogel. There are a variety of materials suitable to load the aerogel medium to produce short-wavelength scattering (e.g. insulating ceramics, high-resistivity silicon, amorphous glass). Diamond was chosen because it has low electromagnetic absorption, high dielectric contrast with respect to vacuum in the microwave through the visible spectral ranges, and because it is commercially available in a wide range of particle sizes [25, 26].
@@ 2.1 Synthesis of polyimide aerogel film embedded with diamond particles
Polyimide synthesis using BPDA as the dianhydride, a combination of diamines, DMBZ and BAPB, and a cross-linking monomer, TAB, iss depicted in Fig. 1. The diamond particle size and loading content are listed in Table LABEL:tab:Aerogel_Fab_run_table in the Supplemental Material. For Runs 1–35, the polymer concentration was formulated at 10 wtand molar ratio of dianhydride to the total diamine fraction was 26:25, with a molar ratio of DMBZ to BAPB at 3:1. For example, in Run 1, 6.94 g BAPB and 23.07 g BPDA were mixed in 320.20 mL NMP and stirred with a magnetic stirrer for 30 min. Then a solution of 12.01 g DMBZ in 20.00 mL NMP solution was added while continuing stirring until the sol was clear and an intermediate amic acid is formed, which is followed by addition of the crosslinker (0.80 g TAB in 10.00 mL NMP). At this stage of synthesis, the diamond particles of different sizes and concentrations were added prior to chemical imidization using 59.30 mL acetic anhydride and 10.93 mL triethylamine. After stirring for 15 min, the sol was poured into molds and aged over 24 hours to ensure full gelation. Molds include 52.5 cm $\times$ 52.5 cm vertical molds with adjusted spacer thickness, a 3-inch diameter round silicone mold, and 20 mL syringe molds with the tips removed. Films were also created by casting the sol on a polyethylene terephthalate (PET) carrier film on a benchtop film casting system, EZ coater EC100, using a doctor blade with the gap set to the desired thickness. The resulting gels were transferred to a solvent bath of 50% NMP + 50% acetone, followed by 100% acetone solvent swap four times per day. The polyimide gels undergo supercritical fluid extraction using liquid CO2 to obtain the final aerogels.
@@ 2.2 Chemical and Mechanical Characterization
where $\rho_{\textrm{b}}$ is the bulk density and $\rho_{\textrm{s}}$ is the skeletal density measured by helium pycnometry. [27] Polyimide aerogel samples were sputter coated with a 5 nm layer thickness of platinum to obtain scanning electron micrographs, using a Hitachi S-4700 field emission scanning microscope. Brunauer-Emmett-Teller (BET) surface area was measured with a Tristar II 3020 Surface Area/Pore Distribution analyzer (Micromeritics Instrument Corp.). Before running nitrogen-adsorption Porosimetry, the samples were out-gassed at 80 C overnight. Thermal conductivities of the samples were measured on a calibrated hot plate device (designed by Aerogel Technologies, LLC) using a flat monolithic sample in direct contact with a plate of NIST reference material of similar planar dimensions with known thermal conductivity. The sample and reference material were sandwiched between an aluminum plate maintained at 37.5 C and a flat-bottomed aluminum vessel filled with an ice bath. Temperatures are measured at the hot side, cold side, and at the interface between the two materials. The heat flux through the material is calculated from the temperature drop across the NIST standard. This heat flux is then used with the thickness and temperature drop across the unknown sample to calculate the sample’s thermal conductivity. Optical microscope images of polyimide aerogel films were captured using a Nikon SMZ 1270 optical microscope.
@@ 2.3 Results and Discussion
The right panel of Fig. 3 shows the measured porosity of the polyimide aerogel with diamonds. As the diamond concentration increases to 80 mg/cm3, the porosity of the aerogel composites slightly decreased from 92% to 89%. This decrease is largely explained by the volume taken up by the diamond particles in the composite, which is approximately 2.3% at 80 mg/cm3, indicating that the aerogel pore structure is not strongly affected by the presence of the diamond particles. The mesoporous/macroporous polyimide network of the aerogels, with typical pore sizes of 5–100 nm, wraps around the much-larger ($>1$ µm) diamond particles as opposed to particle deposition occurring within the pores.
@@
Fig. 5 shows the low magnification scanning electron microscope (SEM) images of the polyimide aerogel composites with different particle sizes and 50 mg/cm3. From the images, diamonds can be seen on the surface of the polyimide aerogel network surface. Polyimide aerogels have 5–100 nm pores, which are smaller than the particle sizes, so it is not possible to see the diamond particles within the pores of the polyimide aerogels. As seen in the higher magnification SEM images (Fig. 5 d, e, and f), the polyimide aerogels show fiber connected networks. The SEM images also verify that the mesoporosity does not change with particle incorporation.
Figure 6: Left: Polyimide aerogel scattering filter prepared as a 1mm thick film with diamond scattering particles of three size ranges, 3–6 µm , 10–20 µm , and 40–60 µm (run 32 as described in Table LABEL:tab:Aerogel_Fab_run_table). Right: Optical microscope image of the same filter.
Formulations containing multiple sizes and concentrations of diamond particles are shown in Table LABEL:tab:Aerogel_Fab_run_table. Combinations of diamonds of 3–6 µm, 10–20 µm, 15–30 µm, or 40–60 µm were incorporated into the aerogels in Runs 30–35 with homogeneous diamond distribution. As seen in the left panel of Fig. 6, a 1 mm thick polyimide aerogel film prepared with three diamond particle sizes, 3–6 µm, 10–20 µm, and 40–60 µm, is flexible enough to be rolled up. The diamond particles are homogeneously distributed throughout the filter, as shown in the optical microscope image in the right panel of Fig. 6. In run 36, the larger diamond particles with 250–350 µm settled at the bottom of the aerogels due to gravity. The bottom polyimide aerogels with larger particles show lower surface area 177.8 m2/g than the top, 309.2 m2/g.
@@ 3 Modeling
Aerogels are highly porous materials with typical pore sizes less than 100 nm. At scales comparable to far-infrared wavelengths, the electromagnetic behavior of the aerogel can be thought of as a homogeneous effective medium and modeled using Maxwell Garnett theory[28, 29, 30, 31]. The base aerogel is modeled as a mixture of polymer and air/vacuum, and a loaded aerogel is modeled as a mixture of the base aerogel with embedded diamond particles. Because the base polyimide polymer has a moderate index of refraction ($n=1.77+0.004i$ at 60 icm[32]) and low filling fraction ($\eta<0.1$), the resulting aerogel has a very low index of refraction ($n\leq 1.1$), it can be modeled as mostly vacuum with some small percentage of plastic polymer. Subsequently, adding dielectric particles to the aerogel can be modeled as a mixture of the base aerogel described above, with higher index of refraction particles homogenously embedded inside. The contrast between the low index of refraction base aerogel and the higher index of refraction from the embedded diamond particles in the composite medium produces Mie scattering. Mie scattering of light leads to rejection of light with a wavelength roughly equal or less than the size of the higher index of refraction particles [24, 33].
The diamond-loaded polyimide aerogel filter behavior is well described by a Mie scattering model. PUREON Microdiamant MONO-ECO monocrystalline diamond powders[34] with size ranges of 1–3 µm, 3–6 µm, 6–12 µm, 10–20 µm, 15–30 µm, 20–40 µm, 40–60 µm, and 250–350 µm were chosen to cover a broad range of particle sizes while also having well-controlled distributions of particles within each range. These sizes of scattering media cover a broad range of potential cutoff frequencies in the far-infrared through millimeter wave regimes.
Particle size distributions are approximately Gaussian, with the end ranges representing the approximate three standard deviation limits for particle diameter. Table 1 shows a sub-sample of particle size distribution data. Diamond powder composition, particle sizes, and shapes were later analyzed with optical microscopy and Raman spectroscopy. Fig. 2 shows images of the diamond particles.
@@
The diamond particle size and loading density are controlled to tune the filter to the desired cutoff frequency. Particle size has the most dominant effect on cutoff frequency. Larger particles have a lower cutoff frequency and smaller particles have a higher cutoff frequency. For a chosen particle size distribution, increasing the loading density creates a filter with a lower cutoff frequency and vice versa. Beyond only changing the cutoff frequency, for a chosen particle loading density and filter thickness, smaller particles reject light more efficiently than larger particles. Geometric scattering cross-section is proportional to the square of the particle radius (area $\propto r^{2}$), while mass density scales with to the cube of particle radius (volume $\propto r^{3}$). Thererore, for a given loading density, larger-radius particles present a smaller total scattering cross-section than smaller-radius particles. Low loading densities of large particles means the particles are on average very far apart and light can “leak” through the material. Low densities of very small particles will still have many particles present to scatter light. Increasing the thickness of the aerogel filters with the largest particles can mitigate these effects.
@@
Figure 7: Left: Demonstration of the effects of scattering particle loading density and radius on modeled aerogel filter cutoff frequency. All three curves show the behavior of a 1 mm thick sample. Peaked features in the yellow model above the cutoff frequency are artifacts of modeling and are not present in measured samples. Right: Modeled and measured specular transmittances near the cutoff for samples D and F as described in Table 2.
In order to improve our modeling abilities, our model can now incorporate multiple scatter size distributions, with different statistical profiles, e.g. Gaussian, uniform, skew, etc. By combining larger particles with smaller particles, we can reduce or eliminate any “blue light leak” above the desired cutoff frequency. Building upon our optical characterization discussed in Section 4 we have incorporated data about the unloaded aerogel performance into the model. The original model estimated the base aerogel matrix dielectric properties without any knowledge of the unique absorption and emission bands of the polymer chains that are not captured by just index of refraction or dielectric constant measurements. Infrared FTS measurements allow us to measure the transmittance, reflectance, and loss of the base aerogel matrix and find any atomic or molecular absorption features and account for them. As Fig. 8 shows, many molecular absorption features appear in the transmittance spectra of unloaded BAPB-based and DADD-based aerogel films between 10 and 100 THz. With this knowledge we can more accurately predict the filter performance and tune filter performance to specific mission requirements as desired. We can also subtract the effects of the base aerogel model to confirm that the behavior of the added scattering particles is Mie-like in reality. With this iterative approach, we are able to closely model the optical behavior of our aerogel scattering filters.
@@ 4.1 Facilities
The PerkinElmer LAMBDA 950 instrument is a dual-beam monochromator that operates at optical wavelengths of 200–3300 nm. We use a photomultiplier tube in the 200–900 nm range and a PbS photocell detector in the 800–3300 nm range. The instrument’s spectral accuracy is <±absentplus-or-minus<$\pm$< ±0.3 nm. The PerkinElmer instrument operates at ambient atmospheric pressure.
@@ 4.2 Aerogel Optical Performance
Because of the improvements made to the Mie Scattering model and the tightly controlled particle size distributions, we are able to design filter recipes to meet specific criteria. There is good agreement between modeled and measured transmittance spectra. Fig. 8 shows transmittance and optical loss measurements of six different aerogel and composite samples. Optical loss is calculated as:
@@
where $\alpha_{\nu}$ is the loss per unit length at frequency $\nu$, $T_{\nu}$ is the specular transmittance at frequency $\nu$ and $d$ is the thickness of the sample. To ensure calculated loss values remain physically-feasible, the lower limit of $T_{\nu}$ is set at $1\times 10^{-5}$ based on the noise floor of our highest-frequency detectors. Although two different polyimide aerogel base formulations were investigated, we chose the BAPB-based formulation described in Section 2 for the remaining scattering filters we characterized optically because the unloaded BAPB sample shows slightly stronger mechanical properties and higher transmittance below 100 THz. The right panel of Fig. 9 shows the measured versus modeled cutoff frequencies for a range of samples. Broadly we find good agreement for all of the samples tested.
Figure 8: Far-infrared specular transmittance spectra and optical loss of unloaded polyimide aerogel material as well as representative samples of aerogel scattering filters. For descriptions of each sample, see Table 2. Thicknesses of measured samples are as follows: A: 0.11 mm, C: 0.18 mm, D: 0.12 mm, E: 0.18 mm, F: 0.1 mm, G: 0.18 mm. (Color scheme from [36].)
Because the range of particles available to us span two orders of magnitude in radius, we are able to construct filters that span a wide range of parameter space. Table 2 shows a selection of different mean particle sizes, loading densities, and subsequent cutoff frequencies available to us. We define the cutoff frequency for a filter as the 50% transmittance point. Larger particles can make cutoff frequencies as low as 2.5 THz, while smaller particles will produce cutoff frequencies at approximately 20 THz.
@@
Figure 11: Measured polyimide aerogel filter density versus index of refraction. Index of refraction for one sample can be determined in two different ways. Index of refraction is for the millimeter and microwave regimes, roughly 500 µm and longer. The two data points for the same sample are connected via a gray line. Data in blue are the calculated indices using the Maxwell Garnett effective medium approximation along with the sample density and the bulk electromagnetic properties of diamond and polyimide. Density of the samples is measured after drying. Data in red are the sample indices determined by fitting data from the VNA to an electromagnetic model of a homogeneous dielectric slab. Due to sample variations in thickness and particle homogeneity, measured VNA data differ slightly from the ideal case considered using the Maxwell Garnett method. Black line shows the expected index for an idealized unloaded polyimide aerogel of varying densities.
@@
A selection of VNA measurement determined sample parameters appear in Table 2. A subset of existing samples was chosen to effectively cover the range of parameter space available. Samples all show indices of refraction of $n\leq 1.08$. Even though VNA measurements were performed from 140–220 GHz ($\simeq$ 1.4–2.1 mm or $\simeq$ 4.7–7.3 icm) the index of refraction is valid across the millimeter and microwave regimes. Because of the expectedly low loss nature of the constituent materials (diamond and polyimide), the samples would need to be $\sim$30 times longer to be able to appropriately measure intrinsic loss with this quasi-optical measurement strategy. Maxwell Garnett theory predicts the loss tangent, $\tan\delta=\frac{\epsilon^{\prime\prime}}{\epsilon^{\prime}}$, to be $\tan\delta\leq 0.001$. Measured in waveguide from 26–40 GHz, the dielectric function of the BAPB-based aerogel without diamond loading is $\epsilon^{*}=1.16+0.0005i$.
@@
The dielectric slab model assumes that samples are perfectly plane parallel, which is not the case for the actual samples as manufactured. By using fitting the VNA data with a dielectric model [37] that incorporates a decoherence factor, $g$, where $g=1$ is perfectly coherent and $g=0$ is perfectly incoherent, we find the samples to have a range of $0.9990.95<g<$ 0.999. Thickness variations, deviations from planarity, (e.g. wedged or curves samples) and structural defects like bubbles will decrease the coherence of the mm-wave light passing through the sample, producing an effect that is degenerate with higher intrinsic material losses. Similarly, the index of refraction from VNA measurements may differ from the homogeneous dielectric slab model because the above effects will alter the coherence of the reflections from each interface. Additionally, during polymerization, the samples typically lay with the optical axis parallel with the gravitational vector. Gravity pulls the diamond particles down to one side and the effect is increased as the particles increase in size. This creates a polymer that has a more densely loaded and a less densely loaded side. The less dense side will have a lower reflectivity than expected and vice versa. In certain applications a gradient index may be desirable when very high loading densities are needed.
Fig. 11 shows a comparison between the index of refraction for the aerogel samples in Table 2 measured in two different ways. The first way is by measuring the densities of the dried samples and then calculate an index from the Maxwell Garnett approximation. The second method uses VNA reflectance data and the fitting parameters from the dielectric slab model, as mentioned above.
The disagreement between the two data sets in Fig. 11 is due to a confluence of factors. All aerogels shrink during the final drying phase, increasing in density slightly from their pre-dried state. Various factors during fabrication can affect the degree to which samples shrink. Shrinkage can be anticipated and taken into account during fabrication, but cannot be fully controlled. Unfortunately there is no way to determine the final density of the diamond loading and the density of the base aerogel separately, so approximate densities are calculated by comparing a loaded dried sample and an unloaded dried sample of the same base formulation.
Regardless of the above effects, prototype filters generally display the desired characteristics of low index and low loss.

### v3

@@
We have developed a suite of infrared-blocking filters made by embedding diamond scattering particles in a polyimide aerogel substrate. We demonstrate the ability to tune the spectral performance of the filters based on both the composition of the base aerogel material and the properties of the scattering particles. We summarize the fabrication, optical modeling, and characterization of these filters. We investigate two polyimide base aerogel formulations and the effects of loading them with diamond scattering particles of varying sizes and relative densities. We describe a model for the filters’ behavior using a combination of Maxwell Garnett and Mie scattering techniques. We present optical characterization results for diamond-loaded aerogel filters with cutoff frequencies (50% transmittance) ranging between 2.5 and 15 THz, and confirm that the measured spectral performance is in agreement with our optical models. We also measure the filters’ refractive indices in the microwave and report findings in agreement with Maxwell Garnett model predictions (typically $n<1.08$).
@@ 1 Introduction
Due to the high porosity of aerogels, they can serve as a host system to other materials without significantly compromising their performance as lightweight, low density, low permittivity materials. The available free volume of the aerogel material allows for the incorporation of particles of higher-index material during fabrication (see Section 2.1). The resulting composite material consists of a network of high-index particles held in suspension by the low-index aerogel[24]. Such particle-loaded polymeric aerogel composites can subsequently be used as optical low-pass filters in the far-infrared through microwave regimes. Incoming light in the passband is transmitted through the low-reflectance and low-loss base medium while unwanted higher-frequency light is scattered and reflected by the composite aerogel. There are a variety of materials suitable to load the aerogel medium to produce short-wavelength scattering (e.g. insulating ceramics, high-resistivity silicon, amorphous glass). Diamond was chosen because it has low electromagnetic absorption, high dielectric contrast with respect to vacuum in the microwave through the visible spectral ranges and is commercially available in a wide range of particle sizes [25, 26].
@@ 2.1 Synthesis of polyimide aerogel film embedded with diamond particles
Polyimide synthesis using BPDA as the dianhydride, a combination of diamines, DMBZ and BAPB, and a cross-linking monomer, TAB, is depicted in Fig. 1. The diamond particle size and loading content used for each formulation are listed in the Supplement 1. For Runs 1–35, the polymer concentration was formulated at 10 wtand molar ratio of dianhydride to the total diamine fraction was 26:25, with a molar ratio of DMBZ to BAPB at 3:1. For example, in Run 1, 6.94 g BAPB and 23.07 g BPDA were mixed in 320.20 mL NMP and stirred with a magnetic stirrer for 30 min. Then a solution of 12.01 g DMBZ in 20.00 mL NMP solution was added while continuing stirring until the sol was clear and an intermediate amic acid is formed, which is followed by addition of the crosslinker (0.80 g TAB in 10.00 mL NMP). At this stage of synthesis, the diamond particles of different sizes and concentrations were added prior to chemical imidization using 59.30 mL acetic anhydride and 10.93 mL triethylamine. After stirring for 15 min, the sol was poured into molds and aged over 24 hours to ensure full gelation. Molds include 52.5 cm $\times$ 52.5 cm vertical molds with adjusted spacer thickness, a 3-inch diameter round silicone mold, and 20 mL syringe molds with the tips removed. Films were also created by casting the sol on a polyethylene terephthalate (PET) carrier film on a benchtop film casting system (ChemInstruments EZ coater EC100) using a doctor blade with the gap set to the desired thickness. The resulting gels were transferred to a solvent bath of 50% NMP + 50% acetone, followed by 100% acetone solvent swap four times per day. The polyimide gels undergo supercritical fluid extraction using liquid CO2 to obtain the final aerogels.
@@ 2.2 Chemical and Mechanical Characterization
where $\rho_{\textrm{b}}$ is the bulk density and $\rho_{\textrm{s}}$ is the skeletal density measured by helium pycnometry [27]. Polyimide aerogel samples were sputter coated with a 5 nm layer thickness of platinum to obtain scanning electron micrographs, using a Hitachi S-4700 field emission scanning microscope. Brunauer-Emmett-Teller (BET) surface area was measured with a Tristar II 3020 Surface Area/Pore Distribution analyzer (Micromeritics Instrument Corp.). Before running nitrogen-adsorption porosimetry, the samples were out-gassed at 80 C overnight. Thermal conductivities of the samples were measured on a calibrated hot plate device (designed by Aerogel Technologies, LLC) using a flat monolithic sample in direct contact with a plate of National Institute of Standards and Technology (NIST) reference material of similar planar dimensions with known thermal conductivity. The sample and reference material were sandwiched between an aluminum plate maintained at 37.5 C and a flat-bottomed aluminum vessel filled with an ice bath. Temperatures were measured at the hot side, cold side, and at the interface between the two materials. The heat flux through the material was calculated from the temperature drop across the NIST standard. This heat flux was then used with the thickness and temperature drop across the unknown sample to calculate the sample’s thermal conductivity. Optical microscope images of polyimide aerogel films were captured using a Nikon SMZ 1270 optical microscope.
@@ 2.3 Results and Discussion
The right panel of Fig. 3 shows the measured porosity of the polyimide aerogel versus diamond concentration. As the diamond concentration increases to 80 mg/cm3, the porosity of the aerogel composites slightly decreases from 92% to 89%. This decrease is largely explained by the volume taken up by the diamond particles in the composite, which is approximately 2.3% at 80 mg/cm3, indicating that the aerogel pore structure is not strongly affected by the presence of the diamond particles. The mesoporous/macroporous polyimide network of the aerogels, with typical pore sizes of 5–100 nm, wraps around the much-larger ($>1$ µm) diamond particles.
@@
Fig. 5 shows scanning electron microscope (SEM) images of three polyimide aerogel composites loaded with diamond particles of different sizes at 50 mg/cm3 loading density. Diamond particles are visible on the surface of the polyimide aerogel network at low magnification (Fig. 5 a, c, and e). Higher magnification SEM images (Fig. 5 b, d, and f) in regions away from diamond particles show the fiber connected networks within the polyimide aerogels. The SEM images also verify that the mesoporosity does not change with particle incorporation.
Figure 6: Left: Polyimide aerogel scattering filter prepared as a 1mm thick film with diamond scattering particles of three size ranges, 3–6 µm , 10–20 µm , and 40–60 µm (run 32 as described in Supplement 1). Right: Optical microscope image of the same filter.
Formulations containing multiple sizes and concentrations of diamond particles are shown in Supplement 1. Combinations of diamonds of 3–6 µm, 10–20 µm, 15–30 µm, or 40–60 µm were incorporated into the aerogels in Runs 30–35 with homogeneous diamond distribution. As seen in the left panel of Fig. 6, a 1 mm thick polyimide aerogel film prepared with three diamond particle sizes, 3–6 µm, 10–20 µm, and 40–60 µm, is flexible enough to be rolled up. The diamond particles are homogeneously distributed throughout the filter, as shown in the optical microscope image in the right panel of Fig. 6. In run 36, the larger diamond particles with 250–350 µm settled at the bottom of the aerogels due to gravity. The bottom regions of these samples show lower surface area 177.8 m2/g than the top, 309.2 m2/g.
@@ 3 Modeling
Aerogels are highly porous materials with typical pore sizes less than 100 nm. The electromagnetic behavior of the aerogel can be thought of as a homogeneous effective medium and modeled using Maxwell Garnett theory[28, 29, 30, 31] at wavelengths longward of the infrared. The base aerogel is modeled as a mixture of polymer and air/vacuum, and a loaded aerogel is modeled as a mixture of the base aerogel with embedded diamond particles. Because the base polyimide polymer has a moderate index of refraction ($n=1.77+0.004i$ at 167 µm [32]) and low volume-filling fraction, the resulting aerogel has a very low index of refraction ($n\leq 1.1$). For the diamond-loaded aerogel filters described in Section 2, the contrast between the low index of refraction base aerogel and the higher index of refraction from the embedded diamond particles in the composite medium produces Mie scattering. Mie scattering of light leads to rejection of light with a wavelength roughly equal or less than the size of the higher index of refraction particles [24, 33].
The diamond-loaded polyimide aerogel filter behavior is well described by a Mie scattering model. PUREON Microdiamant MONO-ECO monocrystalline diamond powders[34] with size ranges of 1–3 µm, 3–6 µm, 6–12 µm, 10–20 µm, 15–30 µm, 20–40 µm, 40–60 µm, and 250–350 µm were chosen to cover a broad range of particle sizes while also having well-controlled distributions of particles within each range. These sizes of scattering media cover a broad range of potential cutoff frequencies in the far-infrared through millimeter wave regimes. Particle size distributions are approximately Gaussian, with the end ranges representing the approximate three standard deviation limits for particle diameter. Table 1 shows a sub-sample of particle size distribution data. Diamond powder composition, particle sizes, and shapes were later analyzed with optical microscopy and Raman spectroscopy. Fig. 2 shows images of the diamond particles.
@@
The diamond particle size and loading density are controlled to tune the filter to the desired cutoff frequency. Particle size has the dominant effect on cutoff frequency. Larger particles lead to lower cutoff frequencies and smaller particles lead to higher cutoff frequencies. For a given particle size distribution, increasing the loading density results in a filter with a lower cutoff frequency and vice versa. Beyond the cutoff frequency, out-of-band blocking performance is determined by particle size, loading density, and filter thickness. In general, smaller particles reject light more efficiently than larger particles. Geometric scattering cross-section is proportional to the square of the particle radius (area $\propto r^{2}$), while mass density scales with to the cube of particle radius (volume $\propto r^{3}$). Therefore, for a given loading density, larger-radius particles present a smaller total scattering cross-section than smaller-radius particles. Low loading densities of large particles means the particles are on average very far apart and light can “leak” through the material. Low densities of very small particles will still have many particles present to scatter light. Increasing the thickness of the aerogel filters with the largest particles can mitigate these effects and improve out-of-band performance.
@@
Figure 7: Left: Demonstration of the effects of scattering particle loading density and radius on modeled aerogel filter cutoff frequency. All three curves show the behavior of a 1 mm thick sample. Peaked features in the orange model above the cutoff frequency are artifacts of modeling and are not present in measured samples. Right: Modeled and measured specular transmittances near the cutoff for samples D and F as described in Table 2.
In order to improve upon the approach in [24], the model has been refined to incorporate multiple scatter size distributions, with different statistical profiles, including Gaussian, uniform, and skew profiles. By combining larger particles with smaller particles, we can reduce or eliminate any “blue light leak” above the desired cutoff frequency. For simplicity, the model presented in [24] used a constant dielectric function for the base aerogel. Building upon our optical characterization discussed in Section 4 we have incorporated the properties of the unloaded aerogel performance into the model. To capture the absorption and emission bands of the polymer chains, spectral measurements of the transmittance and reflectance are captured using the techniques described in Section 4. These data are used to build a dielectric function for the base aerogel matrix across a wide range of infrared frequiencies. As Fig. 8 shows, many molecular absorption features appear in the transmittance spectra of unloaded BAPB-based and DADD-based aerogel films between 10 and 100 THz. T With this knowledge we can more accurately predict the filter performance and tune filter performance to specific mission requirements as desired. We can also subtract the effects of the base aerogel model to confirm that the behavior of the added scattering particles is Mie-like in reality. With this iterative approach, we are able to closely model the optical behavior of our aerogel scattering filters.
@@ 4.1 Facilities
The PerkinElmer LAMBDA 950 instrument is a dual-beam monochromator that operates at optical wavelengths of 200–3300 nm. We use a photomultiplier tube in the 200–900 nm range and a PbS photocell detector in the 800–3300 nm range. The instrument’s spectral accuracy is <±absentplus-or-minus<$\pm$< ±2 nm. The PerkinElmer instrument operates at ambient atmospheric pressure.
@@ 4.2 Aerogel Optical Performance
As a result of our improvements to the Mie scattering model and the tightly-controlled particle size distributions, we are able to design and fabricate filters that span a wide range of parameter space. Table 2 shows a selection of different mean particle sizes, loading densities, and subsequent cutoff frequencies available to us. We define the cutoff frequency for a filter as the 50% transmittance point. Larger particles can set cutoff frequencies as low as 2.5 THz, while smaller particles produce cutoff frequencies at approximately 20 THz. Fig. 8 shows transmittance and optical loss measurements of six different aerogel and composite samples. Optical loss is calculated as:
@@
where $\alpha_{\nu}$ is the loss per unit length at frequency $\nu$, $T_{\nu}$ is the specular transmittance at frequency $\nu$ and $d$ is the thickness of the sample. To ensure calculated loss values remain physically-feasible, the lower limit of $T_{\nu}$ is set at $1\times 10^{-5}$ based on the noise floor of our highest-frequency detectors. Although two different polyimide aerogel base formulations were investigated, we chose the BAPB-based formulation described in Section 2 for the remaining scattering filters we characterized optically because the unloaded BAPB sample shows slightly stronger mechanical properties and higher transmittance below 100 THz. The right panel of Fig. 9 shows the measured versus modeled cutoff frequencies for a range of samples. Broadly, we find good agreement between modeled and measured transmittance spectra (including cutoff frequencies) for all of the samples we studied.
Figure 8: Far-infrared specular transmittance spectra and optical loss of unloaded polyimide aerogel material as well as representative samples of aerogel scattering filters. For descriptions of each sample, see Table 2. Thicknesses of measured samples are as follows: A: 0.11 mm, C: 0.18 mm, D: 0.12 mm, E: 0.18 mm, F: 0.1 mm. (Color scheme from [36].)
@@
Figure 11: Measured polyimide aerogel filter density versus index of refraction. Index of refraction for one sample can be determined in two different ways. Index of refraction is for wavelengths > 500 µm for the millimeter and microwave regimes. The pair of data points for the same sample are connected via a gray line. Data in blue are the calculated indices using the Maxwell Garnett effective medium approximation adopting the measured sample densities. Data in red are the sample indices determined by fitting data from the VNA to an electromagnetic model of a homogeneous dielectric slab. Black line shows the expected index for an unloaded polyimide aerogel across a range of densities. Due to sample variations in thickness and particle homogeneity, measured VNA data differ from the Maxwell Garnett method using the density measurements. Variations in shrinkage from batch-to-batch cause deviations from the idealized aerogel indices and the density-measurement derived indices.
@@
A selection of sample parameters derived from VNA measurements is shown in Table 2. A subset of existing samples was chosen to effectively cover the range of parameter space available. We find indices of refraction of $n\leq 1.08$. Even though VNA measurements were performed from 140–220 GHz ($\simeq$ 1.4–2.1 mm or $\simeq$ 4.7–7.3 icm) the index of refraction is valid across the millimeter and microwave regimes. Because of the expectedly low volume filing fraction of the constituent materials (diamond and polyimide) relative to vacuum, the samples would need to be $\sim$30 times longer to be able to appropriately measure intrinsic loss with this quasi-optical measurement strategy. Maxwell Garnett theory predicts predicts the ratio of the displacement over the conduction current or the loss tangent is to be $\tan\delta\leq 0.001$. Measured in waveguide from 26–40 GHz, the dielectric function of the BAPB-based aerogel without diamond loading is $\epsilon^{*}=1.16\pm 0.01+0.0005i\pm 0.0002i$
@@
The dielectric slab model assumes that samples are perfectly plane parallel, which is not the case for the actual samples as manufactured. By fitting the VNA data with a dielectric model [37] that incorporates a decoherence factor, $g$, where $g=1$ is perfectly coherent and $g=0$ is perfectly incoherent, we find the samples to have a range of $0.9990.95<g<$ 0.999. Thickness variations, deviations from planarity, (e.g. wedged or curves samples) and structural defects like bubbles will decrease the coherence of the mm-wave light passing through the sample, producing an effect that is degenerate with higher intrinsic material losses. Similarly, the index of refraction from quasi-optical VNA measurements may differ from the homogeneous dielectric slab model because the above effects will alter the coherence of the reflections from each interface. Additionally, during polymerization, the samples typically lay with the optical axis parallel with the gravitational vector. Gravity pulls the diamond particles down to one side and the effect is increased as the particles increase in size. This creates a polymer that has a more densely loaded and a less densely loaded side. The less dense side will have a lower reflectivity than expected and vice versa. In certain applications a gradient index may be desirable when very high loading densities are needed.
Fig. 11 shows a comparison between the index of refraction for the aerogel samples in Table 2 measured in two different ways. The first method uses VNA reflectance data and the fitting parameters from the dielectric slab model, as mentioned above. The second uses measured densities of the dried samples to calculate an index from the Maxwell Garnett approximation using knowledge of the bulk material dielectric data. The disagreement between the two data sets in Fig. 11 is due to a confluence of factors affecting both measurement techniques. All aerogels shrink during the final drying phase, increasing in density slightly from their pre-dried state. Various factors during fabrication can affect the degree to which samples shrink. Shrinkage can be anticipated and taken into account during fabrication, but cannot be fully controlled. For diamond-loaded samples, it is impossible to measure the final densities of the polyimide and diamond individually so only total density is known. Measurements of total loaded sample density are combined with knowledge about the shrinkage of unloaded polyimide aerogels to calculate estimated final densities for the diamond particles and the aerogel matrix. Ideally, fabrication of each diamond-loaded sample would be accompanied by an unloaded witness sample to enable more accurate estimation of final diamond and polyimide. Not all loaded aerogels we studied had witness samples, so measured shrinkages from the same formulation from different batches were used. This adds additional systematic error and increasing potential discrepancy between VNA measurement and density measurement derived index of refraction. Regardless of the aforementioned effects, our prototype filters generally display the desired characteristics of low index and low loss.